TheSmartOne (thesmartone):

Does the function has a zero in the given interval? Explain why. f(x) = x^4 - 5 on [-1, 2]

TheSmartOne (thesmartone):

This is a question on IVT Intermediate value theorem

TheSmartOne (thesmartone):

So when x = -1 f(-1) = (-1)^4 - 5 = 1 - 5 = -4 f(-1) = -4 When x = 2 f(2) = (2)^4 - 5 = 8 - 5 = 3 f(2) = 3

TheSmartOne (thesmartone):

So by IVT there will be some x-value between -1 and 2 that will cross the x-axis. Right? o:

OpenStudy (mathmale):

Hello! I'd suggest you experiment a bit. First, sketch y=x^4. Doesn't have to be a fancy sketch. Next, Sketch z=x^4-5 on the same set of axes. Does this "new" function seem to have any roots? From your graph alone, you may be able to determine whether one or both of these roots lies within the interval [-1,2].

OpenStudy (mathmale):

Your approach, using the IMV, is better than mine. Good for you.

OpenStudy (3mar):

Here is your function graphed. You decide whatever has a root (roots) or not.

TheSmartOne (thesmartone):

Yes, it does https://www.desmos.com/calculator/ljdqge2hew

TheSmartOne (thesmartone):

Haha, yes. I'm learning Calc 1 and we just learned about IVT so they want us to apply that in solving it :)

OpenStudy (3mar):

here

TheSmartOne (thesmartone):

Thank you 3mar, but we aren't solving for the zero. We're simply trying to prove that there will be a certain y-value in a given range of x-values. :)

OpenStudy (3mar):

Ok

OpenStudy (3mar):

I just wanted to help

TheSmartOne (thesmartone):

Yes, that's alright :)